### v3

Abstract
The robustness of fault detection algorithms against uncertainty is crucial in the real-world industrial environment. Recently, a new probabilistic design scheme called distributionally robust fault detection (DRFD) has emerged and received immense interest. Despite its robustness against unknown distributions in practice, current DRFD focuses on the overall detectability of all possible faults rather than the detectability of critical faults that are a priori known. Henceforth, a new DRFD trade-off design scheme is put forward in this work by utilizing prior fault information. The key contribution includes a novel distributional robustness metric of detecting a known fault and a new soft distributionally robust chance constraint that ensures robust detectability. Then a new trade-off design scheme of fault detection under unknown probability distributions is proposed, and this offers a flexible balance between the robustness of detecting known critical faults and the overall detectability against all possible faults. To solve the resulting problem, an exact reformulation is derived and a customized solution algorithm is developed, which includes a sequential optimization procedure and an initialization strategy. Finally, case studies on a simulated three-tank system and a real-world battery cell are carried out to showcase the usefulness of our DRFD method.
keywords:
[…]
{highlights}
A distributional robustness metric of detecting a known fault is proposed.
A soft distributionally robust chance constraint is proposed.
A trade-off design scheme of fault detection is proposed and reformulated exactly.
A customized solution algorithm is developed to solve the reformulation.
Case studies demonstrate the usefulness of our method.
1 Introduction
The growing complexity and interconnectivity of modern industrial processes underscore the critical importance of ensuring system safety and reliability. Advanced fault detection (FD) techniques that can effectively handle various types of uncertainty are thus in urgent need. In the past decades, model-based and data-driven FD methods have been extensively investigated and widely applied in a variety of fields. A summary of related methods can be found in Ding (2008, 2014a). In general, current FD methods are built upon the so-called analytical redundancy, and their main steps include residual generation and residual evaluation (Ding, 2008). The aim of residual generation is to produce an informative indicator that captures the unpredictable variations in dynamic systems and ensures sensitivity to faults. Various methods have been developed for the design of residual generators, including observer-based methods (Ding et al., 2019), parity space-based methods (Chow and Willsky, 1984), and subspace identification methods (SIM) (Ding, 2014b; Jin et al., 2023b).
In practical dynamical systems, there exist various types of uncertain disturbances. Because of this, false alarm rate (FAR) and fault detection rate (FDR) are important performance indices in a probabilistic context and the current research focuses on probabilistic FD design that enjoys robustness against uncertainty. For a residual generator, computing its FAR and FDR entails full knowledge of the distribution of stochastic disturbance. However, this is usually unavailable in engineering practices, and the Gaussian assumption has been mostly used as an approximation (Ding, 2014a; Yang and Fang, 2019; Fezai et al., 2020). On this basis, FD via statistical hypothesis testing has been extensively explored, such as the generalized likelihood ratio test (Fezai et al., 2020) and Mann-Whitney test (Yang and Fang, 2019). Unfortunately, the Gaussian assumption itself may not hold true in practice, resulting in unexpected performance loss of FD.
To overcome this limitation, distribution-free methods that consider the worst-case performance against uncertainty in a prescribed compact set have been profoundly investigated (Wang et al., 2020; Hast et al., 2015). Nevertheless, lacking prior knowledge of the distribution may result in over-conservatism. As a hotspot in operations research community, distributionally robust optimization (DRO) offers a systematic and effective solution to address these drawbacks; more importantly, DRO does not require estimating the full probability distribution but introduces an ambiguity set containing all possible probability distributions with similar statistical attributes instead. The two core elements are the shape and the size of the ambiguity set, by which the expected performance under the worst-case distribution is improved. It is a common practice that the former is chosen by decision-makers, such as moment-based constraints (Delage and Ye, 2010; Zymler et al., 2013) and metric-based ball (Mohajerin Esfahani and Kuhn, 2018; Gao and Kleywegt, 2023; Lam, 2019). The advances in DRO have recently inspired the distributionally robust fault detection (DRFD) scheme (Shang et al., 2021; Wan et al., 2021; Shang et al., 2023; Xue et al., 2022). It concurrently yields a residual generator and the alarm threshold by solving an integrated DRO problem while being robust to uncertainty in probability distributions. Thanks to duality theory, the DRFD problem can be reformulated as a deterministic optimization problem that is easier to solve than the primal problem.
In addition to effectively controlling FAR, the detectability for known critical faults, such as most frequently occurring and hazardous faults, is also essential in practice and sometimes needs to be guaranteed. To this aim, a viable option is to integrate prior knowledge into the design phase, thus lifting the detectability of critical faults (Feng et al., 2020; Van de Sand et al., 2021; Yao et al., 2024). However, in the existing DRFD designs (Shang et al., 2021, 2023), only the overall fault detectability of unknown faults is considered, and the information of critical faults is omitted. Hence, it is worth investigating a robust detection scheme considering detectability against both known and unknown faults in a distributionally robust fashion.
To address these issues, we propose in this paper a soft distributionally robust chance constraint (DRCC) to secure the robust detection of known critical faults by leveraging prior fault information. A novel metric is put forward to evaluate the distributional robustness of detecting known critical faults. Then, a new DRFD trade-off design scheme is formalized by maximizing the robustness of both detecting critical faults and the overall detectability against unknown faults. In contrast to existing DRFD in Shang et al. (2021, 2023), our scheme enables utilization of available fault information and results in a flexible balance between detecting known critical faults and detecting other unknown faults. To tackle the induced challenging DRFD design problem, an exact reformulation is derived and a tailored solution algorithm, including an initialization strategy and a sequential optimization procedure, is developed to solve the design problem effectively. The detection performance and efficiency of our method are demonstrated through case studies on a simulated three-tank system and a realistic lithium-ion battery cell. These comprehensive results highlight that there exists a substantial conflict between detecting known critical faults and detecting unknown faults, which can be effectively addressed by our new DRFD design using a flexible and adjustable trade-off.
The rest of this article is organized as follows. Section 2 revisits the basics of residual generation and evaluation for dynamic systems, as well as the design formulation of DRFD. The main results of this work are shown in Section 3, and case studies are presented in Sections 4 and 5. Section 6 concludes this article.
Notations: We use $\mathbb{N}_{i:j}$ to denote the set of non-negative integer indexes $\{i,\cdots,j\}$. $P\succeq 0$ denotes the positive semi-definiteness of a symmetric $P$, $P(:,i:j)$ denotes the submatrix of $P$ from the $i$th column to the $j$th column, $\lVert P\rVert_{F}$ denotes the Frobenius norm. ${\rm tr}(\cdot)$ and ${\rm det}(\cdot)$ are the trace and the determinant operator of a square matrix. For a vector $f$, $\lVert f\rVert^{2}_{P}=f^{\top}Pf$ and $\lVert f\rVert^{2}$ = $f^{\top}f$. $\delta_{\xi}$ represents the Dirac distribution concentrated at $\xi$. For discrete-time signal $x(k)$, the concatenated vector is denoted by $x_{s}(k)=\left[x(k-s)^{\top}\cdots x(k)^{\top}\right]^{\top}\in\mathbb{R}^{(s+1)n_{x}}$ and the block Hankel matrix of $N\in\mathbb{N}_{+}$ interval $X_{k,s}=\left[x_{s}(k-N+1)\cdots x_{s}(k)\right]\in\mathbb{R}^{(s+1)n_{x}\times N}$. The lower triangular block-Toeplitz operator $\mathcal{T}(A,B,C,D)$ is defined as:
$\mathcal{T}(A,B,C,D)=\begin{bmatrix}D&0&0&\cdots&0\\ CB&D&0&\cdots&0\\ CAB&CB&D&\cdots&0\\ \vdots&\ddots&\ddots&\ddots&\vdots\\ CA^{s-1}B&\cdots&CAB&CB&D\end{bmatrix}.$
2 Preliminaries and Problem Formulation
2.1 Residual Generation and Evaluation
Consider a stochastic discrete-time linear time-invariant (LTI) system described by
$\begin{cases}{x}(k+1)=A{x}(k)+B{u}(k)+B_{d}d(k)+B_{f}{f}(k)\\ {y}(k)=C{x}(k)+D{u}(k)+D_{d}d(k)+D_{f}{f}(k)&\end{cases}$ (1)
where ${x}\in\mathbb{R}^{n_{x}}$, ${y}\in\mathbb{R}^{n_{y}}$, ${u}\in\mathbb{R}^{n_{u}}$, $d\in\mathbb{R}^{n_{d}}$, and ${f}\in\mathbb{R}^{n_{f}}$ are process state, measured output, control input, stochastic disturbance, and additive faults, respectively. $A$, $B$, $B_{d}$, $B_{f}$, $C$, $D$, $D_{d}$, and $D_{f}$ are time-invariant matrices having appropriate dimensions. $G_{u}\left(z\right)=C(zI-A)^{-1}B+D$ denotes the transfer function matrix from $u$ to $y$, where $z$ is the time-shift operator. It is assumed that $(C,A)$ is observable. The stable kernel representation (SKR) is a general way to represent the residual generator for LTI systems (Ding, 2008, 2014a). Under the disturbance- and fault-free conditions, the following relation holds:
$\mathcal{K}(z)\begin{bmatrix}u(z)\\ y(z)\end{bmatrix}\equiv 0,~{}\forall u,$
where $\mathcal{K}(z)=\begin{bmatrix}-\hat{N}(z)&\hat{M}(z)\end{bmatrix}$ can be established from the left coprime factorization of the transfer function matrix $G_{u}\left(z\right)=\hat{M}^{-1}(z)\hat{N}(z)$. Motivated by SKR, a general residual generator is expressed in the following form:
$r(z)=\mathcal{K}(z)\begin{bmatrix}u(z)\\ y(z)\end{bmatrix},$
where $r(z)\in\mathbb{R}^{n_{r}}$ is the residual, and its design form in the time domain essentially takes the form of an affine function combination of fault and disturbance:
$r(k)=P[Wd_{s}(k)+Vf_{s}(k)],$ (2)
where $P$ is the projection matrix that parameterizes the residual generator, and shall be deliberately designed to preserve ensuring sensitivity to faults and robustness to disturbances. $\{W,V\}$ are known coefficient matrices, and it is normally assumed that $W$ has full row rank. $\{W,V\}$ can be constructed from system matrices or available past fault-free data, e.g. using the parity space method (Chow and Willsky, 1984) or SIM (Ding, 2014b). Given known state-space matrices, the parity relation of a given order $s\geq n_{x}$ is established as follows:
$y_{s}(k)=\Gamma_{s}x(k-s)+H_{u,s}u_{s}(k)+H_{d,s}d_{s}(k)+H_{f,s}f_{s}(k),$ (3)
where
$\Gamma_{s}=\begin{bmatrix}C^{\top}&(CA)^{\top}&\dots&\left(CA^{s}\right)^{\top}\end{bmatrix}^{\top},$
$H_{u,s}=\mathcal{T}(A,B,C,D),$
$H_{d,s}=\mathcal{T}(A,B_{d},C,D_{d}),$
$H_{f,s}=\mathcal{T}(A,B_{f},C,D_{f}).$
Thus, the residual generator can be designed based on the parity relation (3):
$r(k)=P\cdot\Gamma^{\perp}_{s}\left[y_{s}(k)-H_{u,s}u_{s}(k)\right],$
where $\Gamma^{\perp}_{s}$ is the basis matrix of the left null space of $\Gamma$, satisfying ${\rm rank}(\Gamma^{\perp}_{s})=sn_{y}-n_{x}$ and $\Gamma^{\perp}_{s}\Gamma_{s}=0$.
For systems with unknown state-space matrices, $\{W,V\}$ can be identified from the subspace composed of available historical input-output data
$Z_{\rm p}=\begin{bmatrix}U_{\rm p}\\ Y_{\rm p}\end{bmatrix}=\begin{bmatrix}U_{k-s_{\rm p}-1,s_{\rm p}}\\ Y_{k-s_{\rm p}-1,s_{\rm p}}\end{bmatrix},~{}Z_{\rm f}=\begin{bmatrix}U_{\rm f}\\ Y_{\rm f}\end{bmatrix}=\begin{bmatrix}U_{k,s_{\rm f}}\\ Y_{k,s_{\rm f}}\end{bmatrix}.$
For example, the data-driven SKR can be realized by QR factorization of past and future datasets $\{Z_{\rm p},Z_{\rm f}\}$(Ding, 2014a, b) :
$\begin{bmatrix}Z_{\rm p}\\ U_{f}\\ Y_{f}\end{bmatrix}=\begin{bmatrix}R_{11}&0&0\\ R_{21}&R_{22}&0\\ R_{31}&R_{32}&R_{33}\end{bmatrix}\begin{bmatrix}Q_{1}\\ Q_{2}\\ Q_{3}\end{bmatrix}.$ (4)
By solving the following equation for $\mathcal{K}_{{\rm d},s}$:
$\mathcal{K}_{{\rm d},s}\begin{bmatrix}R_{21}&R_{22}\\ R_{31}&R_{32}\end{bmatrix}=0.$ (5)
Thus, the residual generator can be constructed as:
$r(k)$ $=P\cdot\mathcal{K}_{{\rm d},s}\begin{bmatrix}u_{s}(k)\\ y_{s}(k)\end{bmatrix}.$ (6)
Additionally, one should set $W$ as an identity matrix by taking the residuals $R_{33}Q_{3}$ as samples of $d_{s}(k)$. For sensor bias and actuator failure, $V$ can also be readily specified (Ding, 2014b; Shang et al., 2023). After obtaining the residual, the residual evaluation process follows the alarm logic:
$\begin{cases}J(r)>J_{{\rm th}}\Rightarrow&\text{fault alarm}\\ J(r)\leq J_{{\rm th}}\Rightarrow&\text{no alarm}\end{cases}$
where $J_{\rm th}$ is the alarm threshold, and the Euclidean norm evaluation function $J(r)=\lVert r\rVert^{2}$ is mostly used. Due to uncertainty in $d(k)$, unnecessary alarms could be triggered during fault-free operation, while missed alarms may be present during the occurrence of faults. Thus, FAR and FDR serve as important indicators to evaluate the performance of residual generators.
Definition 1 (FAR and FDR, Ding (2008)).
Given the threshold $J_{\rm th}$, the FAR and FDR for a particular fault ${f}\neq 0$ of an FD design is defined as
${\rm FAR}$ =$=\mathbb{P}\{\lVert r\rVert^{2}>J_{{\rm th}}\mid f=0\},$ (7)
${\rm FDR}$ =$=\mathbb{P}\{\lVert r\rVert^{2}>J_{{\rm th}}\mid f\neq 0\}.$ (8)
2.2 Problem Formulation of DRFD
A critical point of FD design (2) is the choice of $P$, around which numerous solutions have been developed. Among previous works, the unified solution is a conventional approach, which maximizes the ratio of the gains of $f$ relative to $d$ (Wünnenberg and Frank, 1987):
$P^{*}=\arg\min_{P}\frac{\lVert PW\rVert}{\lVert PV\rVert}=\Lambda_{W}^{-1}U_{W}^{\top}.$ (9)
This can be solved by performing singular value decomposition (SVD) of $W=U_{W}\Lambda_{W}V_{W}^{\top}$ (Ding, 2008). After designing $P^{*}$, the Gaussian assumption with $d\sim\mathcal{N}(0,\Sigma_{d})$ is often adopted, and subsequently, $\lVert r\rVert^{2}$ under fault-free conditions is approximately governed by $\chi^{2}$-distribution. Consequently, given a prescribed FAR, the alarm threshold $J_{\rm th}$ can be calibrated to a specific upper quantile of a $\chi^{2}$-distribution (Ding, 2014a; Fezai et al., 2020; Yang and Fang, 2019).
However, the accuracy of the unified solution (9) may not be satisfying in engineering practice. The reason is that (9) is defined using norm-based sensitivity and suffers from over-conservatism. Furthermore, the Gaussian assumption is fragile, leading to performance misspecification when the uncertainties deviate from the Gaussian distribution. To address this issue, DRFD has emerged as an effective scheme (Shang et al., 2021). It regards $\xi=Wd_{s}(k)$ as the source of uncertainty in residual generation, and thus the relation in (2) simplifies to:
$r=P(\xi+Vf),$ (10)
where the dependence on the time index $k$ and the order $s$ is dropped for brevity. In engineering practice, the true distribution $\mathbb{P}_{\xi}$ of the disturbance is not known exactly, and thus using an ambiguity set $\mathcal{D}$ to capture the uncertainty of $\mathbb{P}_{\xi}$ is a common practice in DRO (Shang et al., 2023; Tan et al., 2024). On this basis, the DRFD design scheme aims at maximizing the overall detectability metric under the worst-case FAR constraint within $\mathcal{D}$:
$\max_{P}$ $~{}\rho\left(P\right)$ (DRFD)
${\rm s.t.}$ $~{}\sup_{\mathbb{P}_{\xi}\in\mathcal{D}}\mathbb{P}_{\xi}\{\xi^{\top}P^{\top}P\xi>1\mid f=0\}\leq\alpha$
where $\alpha\in(0,1)$ is an upper bound of FAR that is intuitive to specify for the user. Since there exist an infinite number of pairs $\{P,J_{{\rm th}}\}$ with equivalent detection performance, it is without loss of generality to set the threshold $J_{{\rm th}}=1$. The objective is to maximize an overall fault detectability metric function $\rho(\cdot)$. It evaluates the impact of an unknown profile $f$ on the residual, and can be chosen as the Frobenius norm metric $\rho_{1}(P)=\|PV\|_{F}^{2}=\mathrm{tr}(V^{\top}P^{\top}PV)$ (Ding et al., 2019) or the pseudo-determinant metric $\rho_{2}(P)=\log{\rm pdet}(V^{\top}P^{\top}PV)=\log\det(\Lambda^{\top}U_{1}^{\top}P^{\top}PU_{1}\Lambda)$ (Shang et al., 2021), where $V=U_{1}\Lambda U_{2}^{\top}$ is the compact SVD and $\Lambda$ is diagonal and nonsingular. By solving (DRFD), the optimal solution $P^{*}$ can safely keep the FAR below $\alpha$ under routine fault-free conditions and maximize the impact of unknown fault $f$ on the residual, giving rise to an integrated DRFD design scheme. Central to the DRFD design is the choice of the ambiguity set $\mathcal{D}$, and the Wasserstein ambiguity set has been prevalent recently, which critically relies on the Wasserstein distance for evaluating the discrepancy between two distributions.
Definition 2 (Wasserstein distance, Kantorovich and Rubinshtein (1958)).
Given two distributions $\mathbb{P}$ and $\mathbb{P}^{\prime}$, whose random $\xi$ is over the support set $\Xi\subseteq\mathbb{R}^{n_{\xi}}$,
$d_{\rm W}(\mathbb{P},{\mathbb{P}}^{\prime})=\inf_{\mathbb{Q}\in\mathcal{Q}}\left\{\int_{\Xi^{2}}\rVert\xi-\xi^{\prime}\lVert\mathbb{Q}\left({\rm d}\xi,{\rm d}\xi^{\prime}\right)\right\},$
where $\mathbb{Q}\in\mathcal{Q}$ is composed of all the joint distributions of $\xi$ and $\xi^{\prime}$ with marginal distributions $\mathbb{P}$ and $\mathbb{P}^{\prime}$.
The Wasserstein distance is known to be the minimum transport cost of moving the probability mass from $\mathbb{P}$ to $\mathbb{P}^{\prime}$ in optimal transport theory (Mohajerin Esfahani and Kuhn, 2018). Based on this, the Wasserstein ambiguity set can be defined as follows.
Definition 3 (Wasserstein ambiguity set, Mohajerin Esfahani and Kuhn (2018)).
Given $N$ independent samples $\{\hat{\xi}_{i}\}_{i=1}^{N}$, and Wasserstein radius $\theta$, the Wasserstein ambiguity set is defined as:
$\mathcal{D}_{\rm W}(\theta;N)=\left\{\mathbb{P}\in\mathcal{M}(\Xi)\left|d_{\rm W}(\mathbb{P},\hat{\mathbb{P}}_{N})\leq\theta\right.\right\},$
where $\mathcal{M}(\Xi)$ is the probability space supported on $\Xi$, $\theta$ is the radius of the ambiguity set, and ^$\hat{\mathbb{P}}_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{\hat{\xi}_{i}}$ is the empirical distribution.
The Wasserstein ambiguity set encapsulates a family of distributions that “resemble" the empirical distribution ^$\hat{\mathbb{P}}_{N}$. With the radius $\theta$ appropriately chosen, $\mathcal{D}_{\rm W}(\theta;N)$ is able to capture the true distribution $\mathbb{P}_{\xi}$ with a high probability in a data-driven fashion. Indeed, the uncertainty $\xi$ in (10) can be seen as the primary residual signal under routine fault-free conditions, so the samples $\{\hat{\xi}_{i}\}_{i=1}^{N}$ can be easily collected from an off-the-shelf residual generator to construct $\mathcal{D}_{\rm W}(\theta;N)$ (Shang et al., 2021).
3 Main result
3.1 A New Trade-off Formulation of DRFD
In engineering practice, some critical faults may be a priori known in advance, and their detectability is of direct interest. However, the design scheme in (DRFD) only maximizes the overall detectability against unknown faults but makes no use of known critical fault information. As a result, it may not perform satisfactorily in the presence of such critical faults.
In this work, it is assumed that the profile of a critical fault, e.g. a sensor fault or an actuator fault, is a priori known as $\bar{f}$. The goal is to effectively utilize the directional information of $\bar{f}$ to improve the detectability of DRFD design. First, a soft DRCC is proposed to robustly secure a high FDR of a critical fault $\bar{f}$:
$\begin{split}\mathbb{P}_{\xi}\left\{\lVert P\left(\xi+V\bar{f}\right)\rVert^{2}>1\right\}\geq\beta-{\frac{1}{\eta}}d_{\rm W}(\mathbb{P}_{\xi},\hat{\mathbb{P}}_{N}),~{}~{}~{}~{}~{}~{}~{}&\\ \forall\mathbb{P}_{\xi}\in\mathcal{M}(\Xi),&\end{split}$ (11)
where $\beta\in(0,1)$ denotes a prescribed lower bound of FDR under the empirical distribution ^$\hat{\mathbb{P}}_{\xi}$, and $\eta>0$ is a decision variable. Intuitively, if $\mathbb{P}_{\xi}=\hat{\mathbb{P}}$, a minimal FDR $\beta$ under $\bar{f}$ can be guaranteed by (11). However, when the true distribution $\mathbb{P}_{\xi}$ deviates from $\hat{\mathbb{P}}$, the FDR under $\bar{f}$ may become lower, and the reduction of FDR is at most $1/\eta\cdot d_{\rm W}(\mathbb{P}_{\xi},\hat{\mathbb{P}}_{N})$. Intuitively, the larger $\eta$, the less sensitivity of FDR to the deviation of $\mathbb{P}_{\xi}$. In this sense, $\eta$ yields a novel metric evaluating the distributional robustness of detecting $\bar{f}$ when $\mathbb{P}_{\xi}$ is not exactly known. [^1] Indeed, $1/\eta$ can be interpreted as the fragility in the robust satisficing scheme (Long et al., 2023). To robustly detect the occurrence of a known $\bar{f}$, a large $\eta$ is preferred. As such, we propose to incorporate (11) in the design problem (DRFD) and maximize $\eta$ jointly with the overall detectability $\rho(\cdot)$. This gives rise to a new DRFD trade-off design, which is formalized as follows:
$\max_{P,\eta\geq 0}$ $~{}\rho\left(P\right)+{\gamma}{\eta}$ (12a)
${\rm s.t.}$ $~{}\sup_{\mathbb{P}_{\xi}\in\mathcal{D}_{\rm W}(\theta,N)}\mathbb{P}_{\xi}\left\{\lVert P\xi\rVert^{2}>1\right\}\leq\alpha$ (12b)
$\begin{split}&~{}\mathbb{P}_{\xi}\left\{\lVert P\left(\xi+V\bar{f}\right)\rVert^{2}\leq 1\right\}\leq 1-\beta+{\frac{1}{\eta}}d_{\rm W}(\mathbb{P}_{\xi},\hat{\mathbb{P}}_{N}),\\ &~{}~{}~{}\forall{\mathbb{P}_{\xi}\in\mathcal{M}(\Xi)}\end{split}$ (12c)
Maximizing the distributional robustness metric $\eta$ in (12) encourages a low reduction of detectability of the critical fault when the true distribution deviates from the empirical one $\hat{\mathbb{P}}$. In this sense, the objective (12a) attempts to trade off robust detection of $\bar{f}$ against the overall detectability $\rho(\cdot)$ via a regularization parameter $\gamma\geq 0$. By considering the robustness of detecting a critical fault $\bar{f}$ of specific interest, the soft constraint (12) offers a new trade-off design scheme as compared to the original design (DRFD). Unlike the hard constraint (12b) on FAR, a soft constraint (12c) on FDR is applied to achieve the robustness of detecting $\bar{f}$. It is worth adding that the soft DRCC (12c) is also useful for handling the inexactness of true fault profile $\bar{f}$. By introducing mutiple soft DRCCs into (12), the proposed DRFD scheme can be readily extended to tackle multiple critical faults:
$\max_{P,\eta_{i}}$ $~{}\rho\left({{P}}\right)+\sum^{K}_{i=0}{\gamma_{i}}{\eta_{i}}$
${\rm s.t.}$ $~{}\sup_{\mathbb{P}_{\xi}\in\mathcal{D}_{\rm W}(\theta,N)}\mathbb{P}_{\xi}\left\{\lVert P\xi\rVert^{2}>1\right\}\leq\alpha$
$\begin{split}&~{}\mathbb{P}_{\xi}\left\{\lVert P\left(\xi+V\bar{f}\right)\rVert^{2}\leq 1\right\}\leq 1-\beta_{i}+\frac{1}{\eta_{i}}d_{\rm W}(\mathbb{P}_{\xi},\hat{\mathbb{P}}_{N}),\\ &~{}\qquad\qquad\qquad\qquad\quad{\forall\mathbb{P}_{\xi}\in\mathcal{M}(\Xi)},~{}\forall i\in\mathbb{N}_{1:K}\end{split}$
$~{}\eta_{i}\geq 0,~{}\forall i\in\mathbb{N}_{1:K}$
where $\beta_{i}$ and $\gamma_{i}$ are tolerance FAR levels and weight parameters corresponding to known faults $\bar{f_{i}}$.
3.2 Reformulation and Solution Algorithm
Due to the infinite number of decision variables $\mathbb{P}_{\xi}$ as well as constraints (12b) and (12c), the distributionally robust chance constrained program (12) is challenging to solve. This subsection seeks to derive a tractable counterpart reformulation alongside an effective solution algorithm. To this end, the following two lemmas are first presented.
Lemma 1.
The following optimization problem on an arbitrary non-empty set $\mathcal{S}$, where the decision variable $\mathbb{P}_{\xi}$ is an arbitrary distribution on probability space $\mathcal{M}\left(\Xi\right)$
$\sup_{\mathbb{P}_{\xi}\in\mathcal{M}(\Xi)}\mathbb{P}_{\xi}\{\xi\in\mathcal{S}\}-{\frac{1}{\eta}}d_{\rm W}(\mathbb{P}_{\xi},\hat{\mathbb{P}}_{N})$ (13)
is equivalent to the optimal value of the following problem:
$\begin{split}\inf_{y}~{}&~{}{\frac{1}{N}}\sum^{N}_{i=1}y_{i}\\ {\rm s.t.}&~{}y_{i}\geq 1-{\frac{1}{\eta}}\inf_{\xi\in\mathcal{S}}\lVert\xi-\hat{\xi}_{i}\rVert,~{}\forall i\in\mathbb{N}_{1:N}\\ &~{}y_{i}\geq 0,~{}\forall i\in\mathbb{N}_{1:N}\end{split}$ (14)
Proof.
The proof can be made by adapting that of Theorem 3.1 in Hota et al. (2019) to the problem under study. Define $\kappa={1/\eta}>0$. By definition of the Wasserstein distance, (13) can be rewritten as
$\sup_{\mathbb{P}_{\xi}\in\mathcal{M}(\Xi)}\mathbb{P}_{\xi}\{\xi\in\mathcal{S}\}-\kappa d_{\rm W}(\mathbb{P}_{\xi},\hat{\mathbb{P}}_{N})$ (15)
$=$ $\sup_{\mathbb{Q}\in\mathcal{Q}}~{}\mathbb{E}_{\mathbb{Q}}\left\{\mathbb{1}_{{\rm cl}(\xi\in\mathcal{S})}-\kappa\lVert\xi-\xi^{\prime}\rVert\right\}$
where the ambiguity set $\mathcal{Q}$ is composed of all possible joint distributions of $\xi$ and $\xi^{\prime}$ with marginal distributions $\mathbb{P}$ and $\hat{\mathbb{P}}$:
$\left.\mathcal{Q}=\left\{\mathbb{Q}\in\mathcal{M}(\Xi^{2})\middle|\begin{array}[]{l}({\xi},{\xi^{\prime}})\sim\mathbb{Q}\\ \mathbb{Q}\left({\xi^{\prime}}=\hat{\xi}_{i}\right)=1/N,~{}i\in\mathbb{N}_{1:N}\end{array}\right.\right\}.$
Then the semi-infinite program (15) can be reformulated as
$\sup_{\mathbb{Q}\in\mathcal{Q}}~{}\mathbb{E}_{\mathbb{Q}}\left\{\mathbb{1}_{{\rm cl}(\xi\in\mathcal{S})}-\kappa\lVert\xi-\xi^{\prime}\rVert\right\}$
$=$ $\sup_{\mathbb{Q}\in\mathcal{Q}}\int_{\Xi^{2}}\left(\mathbb{1}_{{\rm cl}(\xi\in\mathcal{S})}-\kappa\lVert\xi-\xi^{\prime}\rVert\right)\mathbb{Q}({\rm d}\xi,{\rm d}\xi^{\prime})$
$=$ $\frac{1}{N}\sum_{i=1}^{N}\sup_{\mathbb{Q}_{i}\in\mathcal{M}(\Xi)}\int_{\Xi}\left[\mathbb{1}_{{\rm cl}(\xi\in\mathcal{S})}-\kappa\lVert\xi-\hat{\xi}_{i}\rVert\right]\mathbb{Q}_{i}({\rm d}\xi)$
$=$ $\frac{1}{N}\sum_{i=1}^{N}\sup_{\xi\in\Xi}\left[\mathbb{1}_{{\rm cl}(\xi\in\mathcal{S})}-\kappa\lVert\xi-\hat{\xi}_{i}\rVert\right]$
where $\{\mathbb{Q}_{i}\}$ are the conditional distributions of $\xi$ given $\xi^{\prime}=\hat{\xi}_{i}$ and each term in the summation can be considered over $\mathcal{S}$ and $\Xi\backslash\mathcal{S}$ respectively as
$\sup_{\xi\in\Xi}\left[\mathbb{1}_{{\rm cl}(\xi\in\mathcal{S})}-\kappa\lVert\xi-\hat{\xi}_{i}\rVert\right]$ (16a)
$=$ $\max\left\{1-\inf_{\xi\in\mathcal{S}}\kappa\lVert{\xi}-{\hat{\xi}_{i}}\rVert,~{}-\inf_{\xi\in\Xi\backslash\mathcal{S}}\kappa\lVert{\xi}-{\hat{\xi}_{i}}\rVert\right\}$ (16b)
$=$ $\max\left\{1-\kappa\inf_{\xi\in\mathcal{S}}\lVert{\xi}-{\hat{\xi}_{i}}\rVert,~{}0\right\},$ (16c)
where (16c) can be derived by distinguishing ${\hat{\xi}_{i}}$ into two cases:
  • When $\hat{\xi}_{i}\in\mathcal{S}$, (16b) equals to $11-\kappa\inf_{\xi\in\mathcal{S}}\lVert{\xi}-{\hat{\xi}_{i}}\rVert=$ 1 due to $-\inf_{\xi\in\Xi\backslash\mathcal{S}}\kappa\lVert{\xi}-{\hat{\xi}_{i}}\rVert\leq 0$ for given $\kappa\geq 0$.
  • When $\hat{\xi}_{i}\in\Xi\backslash\mathcal{S}$, $\inf_{\xi\in\Xi\backslash\mathcal{S}}\kappa\lVert{\xi}-{\hat{\xi}_{i}}\rVert=0$, so (16b) equals to $\max\left\{1-\kappa\inf_{\xi\in\mathcal{S}}\lVert{\xi}-{\hat{\xi}_{i}}\rVert,~{}0\right\}$.
This yields (14) and completes the proof. ∎
Lemma 1 reformulates the DRCC in (12) into semi-infinite constraints. This allows us to further arrive at semi-definite constraints in the following lemma.
Lemma 2.
Given $\mathcal{S}=\{\xi\mid\lVert P\left(\xi+Vf\right)\rVert^{2}\leq 1\}$, the optimal value $\inf_{\xi\in\mathcal{S}}\lVert{\xi}-{\hat{\xi}}\rVert$ is equivalent to solve the following SDP:
$\begin{split}\max_{\pi\geq 0,p,u}&~{}u\\ {\rm s.t.}~{}&~{}\begin{bmatrix}I&-\hat{\xi}\\ -\hat{\xi}^{\top}&{\hat{\xi}}^{\top}\mathbf{\hat{\xi}}-p\end{bmatrix}\\ &\qquad+\pi\begin{bmatrix}P^{\top}P&P^{\top}PVf\\ f^{\top}V^{\top}P^{\top}P&\lVert PVf\rVert^{2}-1\end{bmatrix}\succeq 0\\ &~{}\begin{bmatrix}p&u\\ u&1\end{bmatrix}\succeq 0\end{split}$ (17)
Proof.
The optimal value of (17) is equal to that of the following semi-infinite optimization problem:
$\max_{p,u}~{}$ $u$ (18a)
${\rm s.t.}$ $~{}u^{2}\leq p$ (18b)
$\begin{split}&(\xi-\hat{\xi})^{\top}(\xi-\hat{\xi})\geq p,\\ &\quad\quad\quad\forall\xi~{}{\rm s.t.}~{}\left(\xi+Vf\right)^{\top}P^{\top}P\left(\xi+Vf\right)\leq 1\end{split}$ (18c)
where (18b) essentially describes a second-order cone, which can be reformulated as a linear matrix inequality (LMI):
$\begin{bmatrix}p&u\\ u&1\end{bmatrix}\succeq 0$ (19)
Due to S-procedure (Pólik and Terlaky, 2007), (18c) can also be equivalently translated into the following LMI:
$\begin{bmatrix}I&-\hat{\xi}\\ -\hat{\xi}^{\top}&{\hat{\xi}}^{\top}\mathbf{\hat{\xi}}-p\end{bmatrix}+\pi\begin{bmatrix}P^{\top}P&P^{\top}PVf\\ f^{\top}V^{\top}P^{\top}P&\lVert PVf\rVert^{2}-1\end{bmatrix}\succeq 0$ (20)
Combining (19) and (20) completes the proof.
Now we are ready to address the DRFD problem (12), by establishing the following result.
Theorem 1.
The DRO problem (12) admits the following exact reformulation:
$\max_{\begin{subarray}{c}\bar{P},\eta,\lambda,y,v,\\ \tau,\pi,t,u,p,q\end{subarray}}$ $~{}\rho\left({\bar{P}^{1/2}}\right)+\gamma\eta$ (21a)
${\rm s.t.}$ $~{}\theta+\frac{1}{N}\sum^{N}_{i=1}y_{i}\leq\lambda\alpha$ (21b)
$~{}y_{i}\geq 0,~{}\forall i\in\mathbb{N}_{1:N}$ (21c)
$~{}y_{i}\geq\lambda-t_{i},~{}\forall i\in\mathbb{N}_{1:N}$ (21d)
$~{}\begin{bmatrix}I&-\hat{\xi}_{i}\\ -\hat{\xi}_{i}^{\top}&{\hat{\xi}}_{i}^{\top}\mathbf{\hat{\xi}}_{i}-q_{i}\end{bmatrix}-\tau_{i}\begin{bmatrix}\bar{P}&0\\ 0&-1\end{bmatrix}\succeq 0,~{}\forall i\in\mathbb{N}_{1:N}$ (21e)
$~{}\begin{bmatrix}q_{i}&t_{i}\\ t_{i}&1\end{bmatrix}\succeq 0,~{}i\in\mathbb{N}_{1:N}$ (21f)
$\frac{1}{N}\sum^{N}_{i=1}v_{i}\leq(1-\beta)\eta$ (21g)
$~{}v_{i}\geq 0,~{}\forall i\in\mathbb{N}_{1:N}$ (21h)
$~{}v_{i}\geq\eta-u_{i},~{}\forall i\in\mathbb{N}_{1:N}$ (21i)
$~{}\pi_{i}\begin{bmatrix}\bar{P}&\bar{P}^{\top}Vf\\ f^{\top}V^{\top}\bar{P}&f^{\top}V^{\top}\bar{P}Vf-1\end{bmatrix}$ (21j)
$~{}~{}~{}~{}~{}~{}+\begin{bmatrix}I&-\hat{\xi}_{i}\\ -\hat{\xi}_{i}^{\top}&{\hat{\xi}}_{i}^{\top}\mathbf{\hat{\xi}}_{i}-p_{i}\end{bmatrix}\succeq 0,~{}\forall i\in\mathbb{N}_{1:N}$ (21k)
$~{}\begin{bmatrix}p_{i}&u_{i}\\ u_{i}&1\end{bmatrix}\succeq 0,~{}\forall i\in\mathbb{N}_{1:N}$ (21l)
$~{}\eta\geq 0,~{}\lambda\geq 0,~{}\bar{P}\succeq 0$ (21m)
$~{}\tau_{i}\geq 0,~{}\pi_{i}\geq 0,~{}\forall i\in\mathbb{N}_{1:N}$ (21n)
Proof.
One can obtain the equivalent reformulations (21b)-(21f) to the DRCC (12c) on FDR in virtue of Lemma 1 and 2. Subsequently, the DRCC (12c) on FAR can be reformulated as (21g)-(21l) in a similar spirit. To eliminate as many bilinear terms as possible, one can introduce new decision variables by $\bar{P}=P^{\top}P$, $\lambda:=1/\lambda$, $y_{i}:=y_{i}/\lambda$, and $v_{i}:=v_{i}\eta$ and this yields the proof. ∎
Despite its exactness, the reformulation (21) is NP-hard and intractable due to the presence of the bilinear matrix inequalities (BMI). Thus a customized solution algorithm is necessitated. Notice that when the Lagrange multipliers $\{\tau_{i}\}^{N}_{i=1}$ and $\{\pi_{i}\}^{N}_{i=1}$ are fixed, (22) becomes an SDP with decision variables $\bar{P},\eta,\lambda$ and $\{y_{i},t_{i},q_{i},v_{i},u_{i},p_{i}\}^{N}_{i=1}$. Importantly, SDPs can be modularly solved by interior-point solvers (Toh, 2018) such as MOSEK (ApS, 2024), SDPT3 (Toh et al., 1999) and CSDP (Borchers, 1999). This motivates the development of an efficient solution algorithm based on sequential optimization. As a result, (21) is first decomposed into several subproblems:
$\begin{split}\max&~{}\rho\left({\bar{P}^{1/2}}\right)+\gamma\eta\\ {\rm s.t.}&~{}\mathcal{J}_{1}(\bar{P},\tau)\leq 0,~{}\mathcal{J}_{2}(\bar{P},\eta,\pi)\leq 0\\ &~{}\bar{P}\succeq 0,\eta\geq 0,\tau\geq 0,\pi\geq 0\end{split}$ (22)
where $\mathcal{J}_{1}(\bar{P},\tau)$ is the optimal value of the subproblem:
$\begin{split}\min_{\lambda,y,\tau,t,q}&~{}\theta+\frac{1}{N}\sum^{N}_{i=1}y_{i}-\lambda\alpha\\ {\rm s.t.}~{}~{}&~{}y_{i}\geq 0,~{}y_{i}\geq\lambda-t_{i},~{}\forall i\in\mathbb{N}_{1:N}\\ &~{}\begin{bmatrix}I&-\hat{\xi}_{i}\\ -\hat{\xi}_{i}^{\top}&{\hat{\xi}}_{i}^{\top}\mathbf{\hat{\xi}}_{i}-q_{i}\end{bmatrix}-\tau_{i}\begin{bmatrix}\bar{P}&0\\ 0&-1\end{bmatrix}\succeq 0,\\ &~{}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\forall i\in\mathbb{N}_{1:N}\\ &~{}\begin{bmatrix}q_{i}&t_{i}\\ t_{i}&1\end{bmatrix}\succeq 0,~{}\forall i\in\mathbb{N}_{1:N},~{}\lambda\geq 0\end{split}$ (23)
and $\mathcal{J}_{2}(\bar{P},\eta,\pi)$ is the optimal value of the subproblem:
$\begin{split}\min_{v,\pi,u,p}&~{}\frac{1}{N}\sum^{N}_{i=1}v_{i}-(1-\beta)\eta\\ {\rm s.t.}~{}&~{}v_{i}\geq 0,~{}v_{i}\geq\eta-u_{i},~{}\forall i\in\mathbb{N}_{1:N}\\ &~{}\pi_{i}\begin{bmatrix}\bar{P}&\bar{P}Vf\\ f^{\top}V^{\top}\bar{P}&\lVert Vf\rVert^{2}_{\bar{P}}-1\end{bmatrix}\\ &~{}\qquad\quad~{}+\begin{bmatrix}I&-\hat{\xi}_{i}\\ -\hat{\xi}_{i}^{\top}&{\hat{\xi}}_{i}^{\top}\mathbf{\hat{\xi}}_{i}-p_{i}\end{bmatrix}\succeq 0,~{}\forall i\in\mathbb{N}_{1:N}\\ &~{}\begin{bmatrix}p_{i}&u_{i}\\ u_{i}&1\end{bmatrix}\succeq 0,~{}\forall i\in\mathbb{N}_{1:N}\end{split}$ (24)
Then, the customized solution algorithm based on sequential optimization can be performed and its details are summarized in Algorithm 1. The idea is to iterate between updating the Lagrange multipliers and maximizing the objective. More precisely, the SDPs (23) and (24) are solved to update $\{\tau_{i}\}^{N}_{i=1}$ and $\{\pi_{i}\}^{N}_{i=1}$ while fixing $\bar{P}$ and $\eta$, which enlarges the feasibility region of (22). After updating $\{\tau_{i}\}^{N}_{i=1}$ and $\{\pi_{i}\}^{N}_{i=1}$, they are fixed and the SDP (22) is solved to lift the objective. This iterative process is repeated until convergence.
Algorithm 1 Solution algorithm for reformulation of new DRFD (22)
1:Coefficient matrix $V$, prior fault $\bar{f}$, additive disturbance samples $\{\hat{\xi}_{i}\}_{i=1}^{N}$, Wasserstein radius $\theta$, tolerable FAR $\alpha$ and desired FDR $\beta$.
2:Solve (25) via Algorithm 1 in Shang et al. (2021) using initial $\bar{P}^{(0)}\leftarrow\epsilon I$ with $\epsilon>0$ sufficiently small and obtain $\bar{P}^{*}$.
3:Compute ${\rm FDR}_{\rm emp}\leftarrow\sum_{i=1}^{N}\mathbb{1}_{{\rm cl}(\lVert\xi_{i}+V\bar{f}\rVert_{\bar{P}^{*}}>1)}/{N}$.
4:if $\rm FDR_{\rm emp}<\beta$ then
5:     Return infeasibility.
6:else
7:     Initialize $\bar{P}^{(0)}\leftarrow\bar{P}^{*}$, $\eta^{(0)}\leftarrow 0$ and $n_{\rm ite}\leftarrow 0$.
8:     while Not convergent do
9:         Solve (23) over $\tau$ with $\bar{P}\leftarrow\bar{P}^{(n_{\rm ite})}$ and obtain $\tau^{(n_{\rm ite})}$.
10:         Solve (24) over $\pi$ with $\bar{P}\leftarrow\bar{P}^{(n_{\rm ite})}$ and $\eta\leftarrow\eta^{(n_{\rm ite})}$, and obtain $\pi^{(n_{\rm ite})}$.
11:         Solve (22) with $\tau\leftarrow\tau^{({\rm ite})}$, $\pi\leftarrow\pi^{({\rm ite})}$ and obtain the solution $\bar{{P}}^{(n_{\rm ite}+1)}$ and $\eta^{(n_{\rm ite}+1)}$
12:         $n_{\rm ite}\leftarrow n_{\rm ite}+1$.
13:     end while
14:     Return $\bar{P}$ and $\eta$.
15:end if
3.3 Initialization Strategy and Parameter Tuning
As an approximation solution strategy, Algorithm 1 is essentially a coordinate descent method, which can produce a local optimal solution and does not guarantee global optimality. Therefore, it is crucial to initialize the algorithm with a good feasible solution $\{\bar{P},\eta\}$ to ensure convergence. First, it is recommended that choosing $\eta\geq 0$ be sufficiently small to fulfill (12c) in the best possible way. Second, to initialize $\bar{P}$, we propose to maximize its sensitivity to the direction of $\bar{f}$ in the objective subject to the DRCC of FAR. This can be formulated as the following problem:
$\max_{\bar{P}}$ $~{}\lVert V\bar{f}\rVert_{\bar{P}}^{2}$ (25)
${\rm s.t.}$ $~{}\sup_{\mathbb{P}_{\xi}\in\mathcal{D}}\mathbb{P}_{\xi}\{\lVert\xi\rVert^{2}_{\bar{P}}>1\mid f=0\}\leq\alpha$
Fortunately, (25) is easier to solve than (22) and can be effectively solved by virtue of Algorithm 1 in Shang et al. (2021). However, it may be the case that the induced initial solution $\{\bar{P}^{(0)},\eta^{(0)}\}$ does not satisfy the soft DRCC (12c). A practical strategy is to inspect whether the FDR under the empirical distribution is lower than $\beta$, c.f. Lines 3-5 of Algorithm 1. If so, the design problem (22) is considered to be infeasible and a smaller value of $\beta$ is suggested. Otherwise, the pair $\{\bar{P}^{(0)},\eta^{(0)}\}$ must be a feasible solution to (22) and thus the sequential optimization in Algorithm 1 can be performed.
Before closing this section, a guideline is provided to inform the selection of all hyperparameters $\{\alpha,\theta,\beta,\gamma\}$ for users.
  • $\alpha$ is the maximal tolerance of FAR that shall be preselected according to user preference. In principle, a high value $\alpha$ may result in “alarm flood". Conversely, a low value of $\alpha$ can give rise to poor detectability.
  • $\theta$ is the Wasserstein radius, which regulates the distributional robustness of the constraint on FAR. To calibrate $\theta$, the $K$-fold cross-validation is a viable option, as summarized in Algorithm 2.
  • $\beta$ is the minimal FDR under the nominal distribution ^$\hat{\mathbb{P}}_{N}$ and thus shall be chosen by the user. It should be noted that the FDR is not always kept below $\beta$ when the real distribution $\mathbb{P}_{\xi}$ deviates from ^$\hat{\mathbb{P}}_{N}$. Thus, the value of $\beta$ shall not be too high in order to ensure robustness and avoid infeasibility of (11).
  • $\gamma$ is a trade-off parameter that can be flexibly adjusted to balance between the overall detectability and the robustness of detecting the particular fault $\bar{f}$. A useful tuning heuristic is to inspect the optimal value of $\eta$ a posteriori. If $\eta^{*}$ is too small, then there is a lack of distributional robustness of detecting $\bar{f}$ and thus one shall turn to a larger $\gamma$ or a smaller $\beta$.
Algorithm 2 Cross-validation for selecting the radius $\theta$
1:Number of folds $K$, samples $\{\hat{\xi}_{i}\}_{i=1}^{N}$, grid $\Theta$ of Wasserstein radius $\theta$, and tolerable FAR $\alpha$.
2:Partition the samples $\{\hat{\xi}_{i}\}_{i=1}^{N}$ into $K$ equally sized folds.
3:for Each $\theta\in\Theta$ do
4:     for $k=1:K$ do
5:         Use the $k$th fold of samples as a test dataset, merge the remaining ones to construct $\mathcal{D}_{\rm W}(\bar{\theta};(K-1)N/K)$ and then solve (DRFD).
6:         Compute ${\rm FAR}_{k}$ on the $k$th fold.
7:     end for
8:     Calculate ${\rm FAR}(\bar{\theta})\leftarrow\sum_{k=1}^{K}{\rm FAR}_{k}/K$.
9:end for
10:Find $\theta^{*}\leftarrow\displaystyle\arg\max_{{\theta}\in\Theta}~{}{{\theta}}~{}~{}~{}{\rm s.t.}~{}{\rm FAR}(\bar{\theta})\leq\alpha$.
11:Return $\theta^{*}$.
4 Case Study on A Simulated Three-tank System
First, the efficacy of the proposed DRFD trade-off design scheme is verified on a simulated three-tank system, which has been widely used as a benchmark in the field of FD (Ding, 2008; Shang et al., 2021, 2023; Xue et al., 2022; Ding, 2014a, b; Ding et al., 2019). As shown in Fig. 1, three water tanks are interconnected with pipes, and two pumps feed water to the tanks. The system input $u(k)=\left[u_{1}~{}u_{2}\right]^{\top}\in\mathbb{R}^{2}$ is the flow rates of pumps $1$ and $2$, the system states $x(k)=\left[x_{1}~{}x_{2}~{}x_{3}\right]^{\top}\in\mathbb{R}^{3}$ is the levels of three tanks, and the outputs $y(k)=\left[x_{1}~{}x_{3}\right]^{\top}\in\mathbb{R}^{2}$ is the measurement of the levels of Tanks $1$ and $3$. By linearizing around the operating point $x_{1}=15$cm, $x_{2}=10$cm, $x_{3}=13$cm with a sampling interval $\Delta x=5s$, state-space matrices in (1) are derived as follows (Shang et al., 2021; Ding, 2008):
$A=\begin{bmatrix}0.8945&0.0048&0.1005\\ 0.0048&0.8500&0.0801\\ 0.1005&0.0801&0.8164\end{bmatrix},~{}B=\begin{bmatrix}0.0317&0.0001\\ 0.0001&0.0309\\ 0.0018&0.0014\end{bmatrix}$
$B_{d}=I_{3},~{}B_{f}=\begin{bmatrix}I_{3}&0_{3\times 2}\end{bmatrix},~{}C=\begin{bmatrix}1&0&0\\ 0&0&1\end{bmatrix}$
$D=0,~{}D_{d}=0,~{}D_{f}=\begin{bmatrix}0_{2\times 3}&I_{2}\end{bmatrix}.$
Meanwhile, the disturbance $d$ is governed by a Laplace distribution with mean $\mu_{d}=0$ and covariance $\Sigma_{d}=0.4I$, which no longer follows the conventional Gaussian assumption.
Figure 1: Schematic of the three-tank system.
[…]
Figure 2: Cross-validation results under different Wasserstein radii and overall detectability
Setting different values of $\{\beta,\gamma\}$, trade-off curves between the overall detectability $\rho(\cdot)$ and the distributional robustness index $\eta$ of detecting $\bar{f}$ are derived and displayed in Figs. 3 and 4. It can be seen that for a given $\beta$, the proposed approach enables to flexibly balance between two design objectives by varying $\gamma>0$. With $\gamma$ increasing, the distributional robustness of detecting $\bar{f}$ is enhanced at the cost of lower values of $\rho(\cdot)$. Besides, when $\beta=0.99$, an extremely small $\eta$ is derived, indicating a lack of distributional robustness of detecting $\bar{f}$. This indicates that $\beta$ shall not be chosen too large.
Figure 3: Trade-off between $\rho(\cdot)$ and $\eta$ in NDR1111
Figure 4: Trade-off between $\rho(\cdot)$ and $\eta$ in NDR2222
[…]
Finally, the convergence performance of Algorithm 1 is depicted in Fig. 6, which shows that the convergence condition is met only within four iterations.
Table 1: FAR and FDR Performance on the Simulated Three-Tank Process
DR1111 NDR1111-A NDR1111-B NDR1111-C NDR1111-D NDR1111-E DR2222 NDR2222-A NDR2222-B NDR2222-C NDR2222-D NDR2222-E
FAR (%) 0.27 0.36 0.38 0.37 0.38 0.36 0.19 0.24 0.24 0.26 0.27 0.24
FDR1111 (%) 3.52 96.34 97.05 99.10 99.06 96.22 65.07 96.91 97.01 98.54 98.43 96.90
FDR2222 (%) 96.17 95.92 96.69 75.87 74.95 95.91 97.29 96.15 95.66 85.26 89.03 96.17
Figure 5: FD results on the simulated three-tank system
Figure 6: The objective function iteration of Algorithm 1. (a): NDR1111-C, (b): NDR2222-C
5 Case Study on A Realistic Lithium-ion Battery Cell
Due to the substantial energy density and high efficiency, lithium-ion batteries have found widespread applications in power grids, electric vehicles and portable electronics (Jin et al., 2023a; Liu et al., 2016). Thus, there is a critical need for advanced FD techniques to ensure the safe operation of battery cells (Wang et al., 2024; Xiong et al., 2019). In this section, we perform the case study using real-world data collected from the charging of a lithium-ion cell. Based on Thevenin’s theorem, the equivalent battery circuit model (Xiong et al., 2019; Wang et al., 2024) consists of an ohmic resistance $R_{\rm o}$, a resistor-capacitor pair$\{R_{\rm par},C_{\rm par}\}$ and open-circuit voltage $U_{\rm oc}$, which are shown in Fig. 7. According to Kirchhoff’s law, its dynamic can be written as:
$\begin{cases}\dot{U}_{\rm par}=-\frac{1}{C_{\rm par}R_{\rm par}}U_{\rm par}+\frac{1}{C_{\rm par}}I_{\rm L}\\ \dot{S}=-\frac{1}{Q}I_{\rm L}\\ U_{\rm t}=U_{\rm oc}-U_{\rm par}-I_{\rm L}R_{\rm o}&\end{cases}$
where $U_{\rm par}$, $I_{\rm L}$, $Q$ and $S$ are the voltage of $R_{\rm par}$, the battery current across the battery, the maximum charge capacity and the state-of-charge. $U_{\rm oc}$ is commonly modeled as a polynomial function related to the $S$ (Jin et al., 2023a; Xiong et al., 2019). Thus, the charging process can be represented as a nonlinear system with state $\left[U_{\rm par}~{}S\right]^{\top}$, input $I_{\rm L}$ and output $U_{\rm t}$.
Figure 7: Equivalent circuit model of a single lithium-ion cell
As illustrated in Fig. 8, the real-world experimental platform consists of a temperature tester, a battery tester, a thermostat, and a computer, where all data, including voltage, current, and temperature, are monitored and sampled at a frequency of $1$Hz. More details about the experimental platform setup can be found in Jin et al. (2023a). The input and output of the system are the current across the battery and the terminal voltage.
Figure 8: Battery cell experimental platform
[…]
Figure 9: FD results on the real battery cell charging
Table 2: FAR, FDR and Distributional Robustness of DRFD Designs on Battery Cells
DR1111 NDR1111-A NDR1111-B NDR1111-C DR2222 NDR2222-A NDR2222-B NDR2222-C
FAR (%) 2.00 2.00 1.33 0.00 2.00 0.67 0.00 0.00
FDR1111 (%) 77.33 78.67 84.67 88.00 77.33 87.33 88.00 88.00
FDR2222 (%) 75.33 74.67 83.33 86.67 75.33 86.67 85.33 84.67
$\eta$ / 0.0351 0.1254 0.1728 / 0.1578 0.1740 0.1761
6 Conclusion
In this work, a new DRFD design scheme was proposed that incorporates prior information of critical faults and safely keeps FAR under control without exactly knowing the probabilistic distribution of uncertainty. Built upon a new soft distributionally robust chance constraint, the proposed approach makes it feasible to balance between a robust detection of critical faults and maintaining detectability of unknown faults. Then, an exact reformulation of the proposed new DRFD design problem was derived as a semi-definite program with BMIs. We developed a tailored solution algorithm, including an initialization strategy and a sequential optimization procedure, which helps to effectively tackle the optimization problem by solving SDPs. Finally, comprehensive case studies on a simulated three-tank system and a realistic lithium-ion battery cell showed that our method enables a more balanced performance than generic DRFD designs. It not only maintains good sensitivity against unknown faults but also ensures a robust detection of known critical faults. In future work, it is worth investigating how to relax the dependence on the exact fault information and further address time-varying disturbances by injecting more probabilistic information.
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grant 62373211.